A job interview at a college attracts 100 candidates of which 30% are females and 70% are males. 40% of both female and male candidates have less than 6 months of work experience and the other 60% have at least 6 months of work experience.

If only one of them will be hired, calculate the probability that he/she is a

(i) Male and has less than 6 months of work experience

(ii) Female or has at least 6 months of work experience


1
Expert's answer
2021-05-20T08:19:02-0400

Solution :-

Total condidates = 100

which 30% are females and 70% are males

Number of male=70

number of female=30


according to question,40% of both female and male candidates have less than 6 months of work experience.


so

number of male, less than 6 months of work experience=40% of 70= 28


number of female, less than 6 months of work experience=40% of 30= 12


according to question

other 60% have at least 6 months of work experience.


number of male, at least 6 months of work experience=60% of 70= 42


number of female, at least 6 months of work experience=60% of 30= 18


now If only one of them will be hired,

then


(i) probability (Male and has less than 6 months of work experience)=28/100=7/25


(ii) probability (Female or has at least 6 months of work experience)


=(number of female + number of condidate who have atleast 6 month exp. - number of female who have atleast 6 month exp.)/100


=[30+(42+18)-18]/100

=72/100

=18/25